# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

### v5

@@ D.3 Equation gadget for predicate $x+y+z\equiv 0\pmod{2}$
As a final note, we used AlphaEvolve to find an equation gadget $H=([3+1+n_{\textup{aux}}],E_{u}\cup E_{f},w)$ (see Figure 8(b)) for $P_{0}^{=}$ clauses on $n_{\textup{aux}}=8$ auxiliary vertices. Below we describe $H$ in the same format as Section D.3. The unforced edges in $E_{u}$ are the following.
@@
This gadget achieves $s^{\prime}(H)=c^{\prime}(H)=10$, giving an alternative approach to prove our main result Theorem 4.1. However, this gadget is significantly more complicated than the $P_{1}^{=}$-based equation gadget from Lemma D.4 achieving $s(H)=c(H)=10$, which is why we chose to write our proofs in terms of $s(H)$ and $c(H)$.
@@
Figure 8: Equation gadgets in the $\textup{MWST}_{f}$ instance. Vertices $\{1,2,3\}$ represent variables in the $3\textup{LIN}({2})$ predicate $P^{=}_{0}$. The red edges edges represent the forced edges. The dashed green edges represent the special edges. All edges without an explicitly labeled weight have weight $1$.

### v6

@@ D.3 Equation gadget for predicate $x+y+z\equiv 0\pmod{2}$
As a final note, we used AlphaEvolve to find an equation gadget $H=([3+1+n_{\textup{aux}}],E_{u}\cup E_{f},w)$ (see Figure 8(b)) for $\texttt{P}^{\equiv 0}_{3,2}$ clauses on $n_{\textup{aux}}=8$ auxiliary vertices. Below we describe $H$ in the same format as Appendix D.3. The unforced edges in $E_{u}$ are the following.
@@
This gadget achieves $s^{\prime}(H)=c^{\prime}(H)=10$, giving an alternative approach to prove our main result Theorem 4.1. However, this gadget is significantly more complicated than the $\texttt{P}^{\equiv 1}_{3,2}$-based equation gadget from Lemma D.4 achieving $s(H)=c(H)=10$, which is why we chose to write our proofs in terms of $s(H)$ and $c(H)$.
@@
Figure 8: Equation gadgets in the $\textup{MWST}_{f}$ instance. Vertices $\{1,2,3\}$ represent variables in the $3\textup{LIN}({2})$ equation $\texttt{P}^{\equiv 0}_{3,2}$. The red edges edges represent the forced edges. The dashed green edges represent the special edges. All edges without an explicitly labeled weight have weight $1$.